OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation two is multiplied by -2 and then the equations are added, the result is _____. A.)-3y = 3 B.)3y = 5 C.)5y = 5

OpenStudy (blurbendy):

2x + y = 3 -2x +4y = 2 5y = 5

OpenStudy (anonymous):

Thank you very much. ^^

OpenStudy (blurbendy):

np :)